Suppose you are thinking of purchasing the stock of WHATUPDAWG, Inc.  In addition to the dividend and...

  1. Suppose you are thinking of purchasing the stock of WHATUPDAWG, Inc.  In addition to the dividend and price from year one you expect it to pay a $4.60 dividend in two years.  You believe you can sell the stock for $28.75 at that time.  You require a return of 10% on investments of this risk.  What is the maximum you would be willing to pay?
  2. Suppose you are thinking of purchasing the stock of Super Dawgs, Inc.  In addition to the dividend and price from year one and two you expect it to pay a $5.29 dividend in three years.  You believe you can sell the stock for $33.06 at that time.  You require a return of 10% on investments of this risk.  What is the maximum you would be willing to pay?
Answer #1

1). r = [{Selling Price + Dividend(Year 1) + Dividend(Year 2)} / Purchase Price]1/n - 1

0.10 = [{$28.75 + $4.60 + $4.60} / Purchase Price]1/2 - 1

(1 + 0.10)2 = $37.95 / Purchase Price

Purchase Price = $37.95 / 1.21 = $31.36

2). r = [{Selling Price + Dividend(Year 1) + Dividend(Year 2) + Dividend(Year 3)} / Purchase Price]1/n - 1

0.10 = [{$33.06 + $5.29 + $5.29 + $5.29} / Purchase Price]1/3 - 1

(1 + 0.10)3 = $48.93 / Purchase Price

Purchase Price = $48.93 / 1.331 = $36.76